# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 2 Preliminaries
Let $\Sigma$ be a signature. A subterm dag over $\Sigma$ is a directed, labeled and acyclic multigraph $G=(V,A,\lambda)$, where $V$ is the set of vertices, $A\subseteq V\times V$ is the multiset of directed arcs and $\lambda:V\rightarrow\Sigma$ is the labeling function such that if $outdegree(v)=m$ for some vertex $v\in V$, then $\lambda(v)\in\Sigma_{m}$. The multiset of directed arcs leaving a vertex is ordered. We denote by $v/i$ the $i$th successor of a vertex $v$. The vertices of $G$ can be interpreted as ground terms over $\Sigma$ by the function ^$\widehat{\;}:V\rightarrow W$ which is defined by recursion: for a vertex $v\in V$, we have $\widehat{v}=\lambda(v)$ if $outdegree(v)=0$ and we have $\widehat{v}=\sigma(\widehat{v/1},\ldots,\widehat{v/m})$ if $outdegree(v)=m$ for some $m>0$ and $\lambda(v)=\sigma$. Let $E$ be a GTES over $\Sigma$. We say that $G$ is a subterm dag for a GTES $E$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)$. Let $E$ and $F$ be GTESs over $\Sigma$. We say that $G$ is a subterm dag for the GTES $E$ and GTES $F$ if   $\widehat{\;}$  is a bijection between $V$ and $ST(E)\cup ST(F)$.
@@
Proposition 2.13.
[3] Let $E$ be a GTES over a signature $\Sigma$, and $n=size(E)$. We can construct in $O(n)$ time a subterm dag for the GTES $E$ and the relation $\tau\langle E\rangle$.
@@
Statement 2.14.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. We can construct in $O(n)$ time a subterm dag for the GTESs $E$ and $F$.
@@
[4] Let $E$ be a GTES over $\Sigma$. There exists an algorithm which, given a subterm dag $G=(V,A,\lambda)$ for $E$, and a relation $\tau$ on $V$, produces the congruence closure $\rho$ of $\tau$ on $V$ in $O(n\,\mathrm{log}\,n)$ time, where $n=size(E)$. The algorithm gives each $\rho$-class a unique name in the set { $\{\,1,\ldots,|V/\rho|\,\}$, and also outputs an array of the cardinalities of all $\rho$-classes.
@@
3 Running Examples
We now start our eight running examples. Throughout this section, $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
@@ 3 Running Examples
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f,\;g\,\}$, $E=\{\,f(\#)\doteq g(\#)\,\}$, and $F=\{\,f(\$)\doteq g(\$)\,\}$. Observe the following equations:
@@
$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\delta(s),\delta(t))\mid\delta\in CON_{\Sigma,1},s,t\in\{\,f(\$),\;g(\$)\,\}\,\}$,
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,f\,\}$, $E=\{\,f(f(\#))\doteq\#\,\}$, and $F=\{\,f(f(f(\#)))\doteq\#\,\}$.
@@
$\mathop{\leftrightarrow}^{*}_{E}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }2\,\}$,
$\mathop{\leftrightarrow}^{*}_{F}=\{\,f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }3\,\}$,
@@
The equivalence classes of $\Theta\langle E,F\rangle$ are {#$\{\,\#,\;f^{2}(\#)\}$, $\{\,f(\#),\;f^{3}(\#)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f^{3}(\#)\}$, $\{\,f(\#)\,\}$,   $\{\,f^{2}(\#)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#$\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$.
@@
#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle F,E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
Consequently,
for each $t\in W$, $\left(t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle E,F\rangle}\mbox{ and }t/_{\Theta\langle{E\cup F},\emptyset\rangle}\neq t/_{\Theta\langle{F},E\rangle}\right)$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\,\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$.
Intuitively, we now define GTESs $E$ and $F$ such that Statements 1 – 4 hold.
@@ 1. $\mathop{\leftrightarrow}^{*}_{E}$ has four congruence classes:
  • the set of all ground terms which are different from the constant ##\##, and contain only the symbols $f$ and ##\##,
  • the set of all ground terms containing only the symbols $f$ and \$currency-dollar\\$\$, and
@@
2. We obtain GTES $F$ from GTES $E$ by swapping the constants ##\## and \$currency-dollar\\$\$.
@@ 3. $\mathop{\leftrightarrow}^{*}_{F}$ has four congruence classes:
  • the set of all ground terms which are different from the constant \$currency-dollar\\$\$, and contain only the symbols $f$ and \$currency-dollar\\$\$,
  • the set of all ground terms containing only the symbols $f$ and ##\##, and
@@ 4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
  • the set of all ground terms which contain only the symbols $f$ and \$currency-dollar\\$\$,
  • the set of all ground terms containing only the symbols $f$ and ##\##, and
@@
$E=\{\,f(\#,f(\#,\#))\doteq f(\#,\#),\;f(f(\#,\#),\#)\doteq f(\#,\#),\;f(f(\#,\#),f(\#,\#))\doteq f(\#,\#),\;\newline f(\$,\$)\doteq\$,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\$,f(\#,\#))\doteq\mathsterling,\;f(\mathsterling,f(\#,\#))\doteq\mathsterling,\;\newline f(f(\#,\#),\$)\doteq\mathsterling,\;f(f(\#,\#),\mathsterling)\doteq\mathsterling\,\}$
$F=\{\,f(\$,f(\$,\$))\doteq f(\$,\$),\;f(f(\$,\$),\$)\doteq f(\$,\$),\;f(f(\$,\$),f(\$,\$))\doteq f(\$,\$),\;\newline f(\#,\#)\doteq\#,\;\newline f(\$,\#)\doteq\mathsterling,\;f(\$,\mathsterling)\doteq\mathsterling,\;\newline f(\#,\$)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\$)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling\newline f(\#,f(\$,\$))\doteq\mathsterling,\;f(\mathsterling,f(\$,\$))\doteq\mathsterling,\;\newline f(f(\$,\$),\#)\doteq\mathsterling,\;f(f(\$,\$),\mathsterling)\doteq\mathsterling\,\}$.
GTES $E$ is symmetric to GTES $F$. Note the following equations:
  • $ST\langle E\rangle=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;\newline f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling)\,\}$,
  • $ST\langle F\rangle=\{\,\$,\;\#,\;\mathsterling,\;\newline f(\$,\$),\;f(\$,\#),\;f(\$,\mathsterling),\;\newline f(\#,\$),\;f(\#,\#),\;f(\#,\mathsterling),\;\newline f(\mathsterling,\$),\;f(\mathsterling,\#),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;\newline f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$,
  • $W=\{\,\#,\;\$,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\$,f(\$,\$)),\;\newline f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$)),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
The equivalence classes of $\Theta\langle E,F\rangle$ are
{#},{\$$\{\,\#\,\},\;\newline \{\,\$,\;f(\$,\$)\,\},\;\newline \{\,f(\#,\#),\;f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;f(f(\#,\#),f(\#,\#))\,\},\newline \;\{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are
{#$\{\,\#,\;f(\#,\#)\,\},\;\newline \{\,\$\,\},\newline \{\,f(\$,\$),\;f(\$,f(\$,\$)),\;f(f(\$,\$),\$),\;f(f(\$,\$),f(\$,\$))\},\;\newline \{\,\mathsterling,\;f(\#,\$),\;f(\#,\mathsterling),\;f(\$,\#),\;f(\$,\mathsterling),\;f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;\newline f(\$,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;f(f(\#,\#),\$),\;f(f(\#,\#),\mathsterling),\;\newline f(\#,f(\$,\$)),\;f(\mathsterling,f(\$,\$)),\;f(f(\$,\$),\#),\;f(f(\$,\$),\mathsterling)\,\}$.
@@
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\#,\#)\,\}\cup(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\times(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}})\cup T_{\{\,\$,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
@@
$\mathop{\leftrightarrow}^{*}_{F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}}))\right)\cup\newline \{\,(\$,\$)\,\}\cup\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)\times\left(T_{\{\,\$,f\,\}}\setminus\{\,\$\,\}\right)$.
@@
$\mathop{\leftrightarrow}^{*}_{E\cup F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\times\left((T_{\Sigma}\setminus(T_{\{\,\$,f\,\}}\cup T_{\{\,\#,f\,\}})\right)\cup\newline (T_{\{\,\$,f\,\}}\times T_{\{\,\$,f\,\}})$.
$\mathop{\leftrightarrow}^{*}_{E\cup F}$ has the congruence classes $T_{\{\,\#,f\,\}}$,   $T_{\{\,\$,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\$,f\,\}})$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,\#\doteq\$,\;\#\doteq\mathsterling,\;\#\doteq f(\#,\#),\;\#\doteq f(\#,\flat),\;\#\doteq f(\flat,\#)\,\}$,
@@
$F=\{\,\mathsterling\doteq\flat,\newline \mathsterling\doteq f(\#,\#),\;\mathsterling\doteq f(\#,\$),\;\mathsterling\doteq f(\#,\mathsterling),\;\newline \mathsterling\doteq f(\$,\#),\;\mathsterling\doteq f(\$,\$),\;\mathsterling\doteq f(\$,\mathsterling),\;\newline \mathsterling\doteq f(\mathsterling,\#),\;\mathsterling\doteq f(\mathsterling,\$),\;\mathsterling\doteq f(\mathsterling,\mathsterling)\,\}$.
@@
$\mathop{\leftrightarrow}^{*}_{F}=\{\,(\#,\#),\;(\$,\$)\,\}\cup((T_{\Sigma}\setminus\{\,\#,\;\$\,\})\times(T_{\Sigma}\setminus\{\,\#,\;\$\,\}))$, and
@@
Note that (#$(\#,\flat)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Consequently,
@@
$W=\{\,\#,\;\$,\;\mathsterling,\;\flat,\;\newline f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\mathsterling),\;f(\mathsterling,\flat),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\}$.
@@
{#,\$,£$\{\,\#,\;\$,\;\mathsterling,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\;\newline \{\,\flat\,\}$.

### v2

@@ 2 Preliminaries
[…]
@@
Proposition 2.14 ([4, Proposition 3.4]).
Let $E$ be a GTES over a signature $\Sigma$, and $n=\mathrm{size}(E)$. We can construct in $O(n)$ time a subterm dag for the GTES $E$ and the relation $\tau\langle E\rangle$.
@@
Statement 2.15.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. We can construct in $O(n)$ time a subterm dag for the GTESs $E$ and $F$.
@@
[9] Let $E$ be a GTES over $\Sigma$. There exists an algorithm which, given a subterm dag $G=(V,A,\lambda)$ for $E$, and a relation $\tau$ on $V$, produces the congruence closure $\rho$ of $\tau$ on $V$ in $O(n\,\mathrm{log}\,n)$ time, where $n=\mathrm{size}(E)$. The algorithm gives each $\rho$-class a unique name in the set { $\{\,1,\ldots,|V/\rho|\,\}$, and also outputs an array of the cardinalities of all $\rho$-classes.
@@
3 Running Examples for the Main Cases
Along the proof of our main result, we shall distinguish four main cases. We define these four main cases, and then we present eight running examples to illustrate our concepts and results. In each main case, a running example presents GTESs $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ and a running example presents $E$ and $F$ such that $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. From now on throughout the paper, $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
  • Main Case 1:
    $\Sigma$ is a unary signature. Example 3.18 and its sequels, Examples 5.51, 7.72, 10.105. and 12.119, and 12.124, also Example 3.19 and its sequels, 5.52, 7.73, 10.106, 12.120, and 12.125 illustrate this case.
  • Main Case 2:
    Both GTRS $R\langle E,F\rangle$ and GTRS $R\langle F,E\rangle$ are total. Example 3.20 and its sequels Examples 5.53, 13.127, further Example 3.21 and its sequels, Examples 5.54, 13.128, exemplify this case.
  • Main Case 3:
    $\Sigma_{k}\neq\emptyset$ for some $k\geq 2$, and at least one of $R\langle E,F\rangle$ and $R\langle F,E\rangle$ is total. Example 3.22 and its sequels, Examples 5.55, 7.74, 10.110, 14.135, and 14.142, moreover Example 3.23, and its sequels, Examples 5.56, 7.75, 10.111, 14.136, and 14.143 present this case.
  • Main Case 4:
For each running example, we compute the sets $ST\langle E\rangle$, $ST\langle F\rangle$, $W$ and the relations $\mathop{\leftrightarrow}^{*}_{E}$, $\mathop{\leftrightarrow}^{*}_{F}$, $\mathop{\leftrightarrow}^{*}_{E\cup F}$, $\Theta\langle E,F\rangle$, $\Theta\langle F,E\rangle$, $\Theta\langle E\cup F,\emptyset\rangle$, and state whether $\mathop{\leftrightarrow}^{*}_{E\cup F}\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$ or $\mathop{\leftrightarrow}^{*}_{E\cup F}\not\subseteq\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$.
Recall that $E$ and $F$ are GTESs over a signature $\Sigma$, and $W$ stands for $ST\langle E\cup F\rangle$.
@@ 3 Running Examples for the Main Cases
[…]
@@
[…]
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{1}=\{\,f\,\}$, $E=\{\,f(f(\#))\doteq\#\,\}$, and $F=\{\,f(f(f(\#)))\doteq\#\,\}$.
@@
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }2\,\}$,
$\mathop{\leftrightarrow}^{*}_{F}=\{\,(f^{m}(\#),f^{n}(\#))\mid m,n\in{\mathbb{N}},\;m-n\mbox{ is divisible by }3\,\}$,
@@
The equivalence classes of $\Theta\langle E,F\rangle$ are $\{\,\#,\;f^{2}(\#)\}$, $\{\,f(\#),\;f^{3}(\#)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are $\{\,\#,\;f^{3}(\#)\}$, $\{\,f(\#)\,\}$,   $\{\,f^{2}(\#)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is $\{\,\#,\;f(\#),\;f^{2}(\#),\;f^{3}(\#)\,\}$.
@@
for each $t\in W$, $\left(t/_{\Theta\langle E,F\rangle}\subset t/_{\Theta\langle{E\cup F},\emptyset\rangle}\mbox{ and }t/_{\Theta\langle{F},E\rangle}\subset t/_{\Theta\langle{E\cup F},\emptyset\rangle}\right)$.
@@
[…]
@@ 1. $\mathop{\leftrightarrow}^{*}_{E}$ has four congruence classes:
  • 1.
    The singleton $\{\,\#\,\}$. Because there is no equation in $E$ with left-hand side or right-hand side #\#. We consider #\# as the representative of this class.
  • 2.
    The set of all ground terms which are different from the constant #\#, and contain only the symbols $f$ and #\#. Applying the first three equations, $E$ rewrites each ground term $t$ of height $2$ in the above set to f(#,#)f(\#,\#). Hence for each ground term $t$ in this set, we have $t\mathop{\leftrightarrow}^{*}_{E}f(\#,\#)$. We consider f(#,#)f(\#,\#) as the representative of this class.
  • 3.
@@
  • 4.
@@
[…]
@@ 3. $\mathop{\leftrightarrow}^{*}_{F}$ has four congruence classes:
  • 2.
[…]
  • 3.
    the set of all ground terms containing only the symbols $f$ and #\#, and
  • 4.
@@ 4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
  • 1.
[…]
  • 2.
    the set of all ground terms containing only the symbols $f$ and #\#, and
  • 3.
@@
$E=\{\,f(\#,f(\#,\#))\doteq f(\#,\#),\;f(f(\#,\#),\#)\doteq f(\#,\#),\;f(f(\#,\#),f(\#,\#))\doteq f(\#,\#),\;\newline f(\mathdollar,\mathdollar)\doteq\mathdollar,\;\newline f(\#,\mathdollar)\doteq\mathsterling,\;f(\#,\mathsterling)\doteq\mathsterling,\;\newline f(\mathdollar,\#)\doteq\mathsterling,\;f(\mathdollar,\mathsterling)\doteq\mathsterling,\;\newline f(\mathsterling,\#)\doteq\mathsterling,\;f(\mathsterling,\mathdollar)\doteq\mathsterling,\;f(\mathsterling,\mathsterling)\doteq\mathsterling,\newline f(\mathdollar,f(\#,\#))\doteq\mathsterling,\;f(\mathsterling,f(\#,\#))\doteq\mathsterling,\;\newline f(f(\#,\#),\mathdollar)\doteq\mathsterling,\;f(f(\#,\#),\mathsterling)\doteq\mathsterling\,\}$,
[…]
Note the following equations:
  • 1.
    $ST\langle E\rangle=\{\,\#,\;\mathdollar,\;\mathsterling,\;\newline f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\mathsterling),\;\newline f(\#,f(\#,\#)),\;f(f(\#,\#),\#),\;\newline f(f(\#,\#),f(\#,\#)),\;f(\mathdollar,f(\#,\#)),\;f(\mathsterling,f(\#,\#)),\;\newline f(f(\#,\#),\mathdollar),\;f(f(\#,\#),\mathsterling)\,\}$,
  • 2.
[…]
  • 3.
[…]
@@
$\mathop{\leftrightarrow}^{*}_{E}=\{\,(\#,\#)\,\}\cup(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\times(T_{\{\,\#,f\,\}}\setminus\{\,\#\,\})\cup(T_{\{\,\mathdollar,f\,\}}\times T_{\{\,\mathdollar,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)\times\left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}})\cup T_{\{\,\mathdollar,f\,\}})\right)$.
$\mathop{\leftrightarrow}^{*}_{E}$ has the congruence classes $\{\,\#\,\}$,   $T_{\{\,\#,f\,\}}\setminus\{\,\#\,\}$,   $T_{\{\,\mathdollar,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})$.
Then
  • 1.
    $T_{\{\,\#,f\,\}}$ is the set of all ground terms containing only the symbols $f$ and #\#.
  • 2.
[…]
  • 3.
[…]
  • 4.
@@
$\mathop{\leftrightarrow}^{*}_{E\cup F}=(T_{\{\,\#,f\,\}})\times T_{\{\,\#,f\,\}})\cup(T_{\{\,\mathdollar,f\,\}}\times T_{\{\,\mathdollar,f\,\}})\cup\newline \left(T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)\times\left((T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})\right)$.
$\mathop{\leftrightarrow}^{*}_{E\cup F}$ has the congruence classes $T_{\{\,\#,f\,\}}$,   $T_{\{\,\mathdollar,f\,\}}$,   and $T_{\Sigma}\setminus(T_{\{\,\#,f\,\}}\cup T_{\{\,\mathdollar,f\,\}})$.
@@
The equivalence classes of $\Theta\langle E,F\rangle$ are
[…]
The equivalence classes of $\Theta\langle F,E\rangle$ are
[…]
@@
[…]
$E=\{\,\#\doteq\mathdollar,\;\#\doteq\mathsterling,\;\#\doteq f(\#,\#),\;\#\doteq f(\#,\flat),\;\#\doteq f(\flat,\#)\,\}$,
@@
$F=\{\,\mathsterling\doteq\flat,\newline \mathsterling\doteq f(\#,\#),\;\mathsterling\doteq f(\#,\mathdollar),\;\mathsterling\doteq f(\#,\mathsterling),\;\newline \mathsterling\doteq f(\mathdollar,\#),\;\mathsterling\doteq f(\mathdollar,\mathdollar),\;\mathsterling\doteq f(\mathdollar,\mathsterling),\;\newline \mathsterling\doteq f(\mathsterling,\#),\;\mathsterling\doteq f(\mathsterling,\mathdollar),\;\mathsterling\doteq f(\mathsterling,\mathsterling)\,\}$.
@@
$)\mathop{\leftrightarrow}^{*}_{F}=\{\,(\#,\#),\;(\mathdollar,\mathdollar)\,\}\cup((T_{\Sigma}\setminus\{\,\#,\;\mathdollar\,\})\times(T_{\Sigma}\setminus\{\,\#,\;\mathdollar\,\}$)), and
@@
Note that $(\#,\flat)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$. Consequently,
@@
$W=\{\,\#,\;\mathdollar,\;\mathsterling,\;\flat,\;\newline f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\mathsterling),\;f(\mathsterling,\flat),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\}$.
@@
$\{\,\#,\;\mathdollar,\;\mathsterling,\;\newline \,f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\;\newline \{\,\flat\,\}$.